November 1996  ICTCM IX (Reno)  0

Jerks Don’t Jump — Mathematical Modelling with an Impact

Douglas B. Meade

Department of Mathematics University of South Carolina Columbia, SC 29208

Phone:  (803) 777-6183

E-mail:  meade@math.sc.edu
WWW URL: http://www.math.sc.edu/˜ meade/

7 November 1996